he energy input to an engine is 3.00 times greater than

the work it performs. What is its thermal efficiency?


1
Expert's answer
2021-06-11T11:22:07-0400
"{\rm Efficiency =\frac{output\: work}{input\: energy}}""=\frac{W}{3W}=\frac{1}{3}=33.3\%"